OpenStudy (anonymous):

Tennis balls are often sold in a can of three. Assuming that the balls are packed tightly so that they touch the lateral side and the bases, determine the amount of volume in the can that is not taken up by the tennis balls

OpenStudy (anonymous):

tennis balls volume is 137.26cm^3

OpenStudy (anonymous):

Does 3.2 cm represent the radius? Find the volume of one tennis ball. V = (4pi/3)r^3 = (4pi/3)*(1.6)^3 = 137.26 cm^3. 3 tennis balls will take up 3 times that volume, so 411.775 cm^3. Then, find the volume of the can, which will have a radius of 3.2 (width of a tennis ball) and a height of 19.2 cm (diameters of 3 tennis balls). The volume of the can is V = pi r^2 * h = 617.662 cm^3. Subtract the 411.775 from this value to get the answer: 205.887 cm^3 not taken by the tennis balls.

OpenStudy (anonymous):

thank you